OpenStudy (31356):

Solve by substitution with 3 variables (Please show your steps) MEDAL AWARDED: x=4 x+y=-6 4x-3y+2z=26

Vocaloid (vocaloid):

hint: x = 4, we can replace "x" with "4" whenever we see it so if x + y = -6, what is y?

OpenStudy (31356):

y=-10

Vocaloid (vocaloid):

good! now we have x = 4 and y = -10 now we can plug those in to the third equation to find z

OpenStudy (31356):

Right...

Vocaloid (vocaloid):

4x-3y+2z=26 16 + 30 + 2z = 26 z = ?

OpenStudy (31356):

z=-10

Vocaloid (vocaloid):

hm, not quite

OpenStudy (31356):

Let me double check...

Vocaloid (vocaloid):

2z = -10 z = ?

OpenStudy (31356):

-5

Vocaloid (vocaloid):

good! so our final answer is (4,-10,-5)

OpenStudy (31356):

Sorry about that, I misread it and solved too fast..

OpenStudy (31356):

Thank you! You make this problem seem so easy!